Given the following piecewise function, evaluate f(-5). I < -4 f(x) = . 1-42 -3x (x² – 2 -4 < x < 0 0<x
Answer #1

Solution given function; f(x) = 5-4x+1 1-3X (²2 rc-4. -45x20 OLX since x=-5<.-4, required function is f(1) = :-4x+1 f(-5) = -